# Do LLMs write like humans? Variation in grammatical and rhetorical styles

### v1

@@ Abstract
Large language models (LLMs) are capable of writing grammatical text that follows instructions, answers questions, and solves problems. As they have advanced, it has become difficult to distinguish their output from human-written text. While past research has found some differences in surface features such as word choice and punctuation, and developed classifiers to detect LLM output, none has studied the rhetorical styles of LLMs.
@@
Using several variants of Llama 3 and GPT-4o, we construct two parallel corpora of human- and LLM-written texts from common prompts. Using Douglas Biber’s set of lexical, grammatical, and rhetorical features, we identify systematic differences between LLMs and humans and between different LLMs. These differences persist when moving from smaller models to larger ones, and are larger for instruction-tuned models than base models. This demonstrates that despite their advanced abilities, LLMs struggle to match human styles, and hence more advanced linguistic features can detect patterns in their behavior not previously recognized.
1 Introduction
@@
As large language models (LLMs) have advanced in recent years, from “stochastic parrots” to models evidently capable of performing complex tasks, most attention has focused on their reasoning performance: solving mathematical problems, writing code, evaluating arguments, diagnosing diseases, and so on (Wang et al., 2023; Huang and Chang, 2023). While past research has studied their mastery of basic grammar and vocabulary (Chang and Bergen, 2024), there is relatively little research on their language performance more generally: their ability to produce readable text in a variety of styles. Rather than exploring it in detail, commentators discuss the business and communication tasks that might be automated by LLMs with their writing ability, or consider the dangers of impersonation and misinformation facilitated by LLMs (Barman et al., 2024; Kumar et al., 2023; Kovács, 2024). As more and more writing tasks are automated, such problems appear inevitable.
@@
However, the impression that LLMs write “like humans” is based primarily on qualitative evaluation of their output, not on thorough linguistic evaluation of their text. So far, quantitative comparisons have looked mainly at basic grammar and syntax (Chang and Bergen, 2024) or surface features such as word choice, punctuation, sentence length, and so on, finding evidence of some differences between human- and LLM-written text (Tang et al., 2024; Fröhling and Zubiaga, 2021; Muñoz-Ortiz et al., 2024; Herbold et al., 2023; Liang et al., 2024). Other work has used these features, or language models trained on sample texts, to classify LLM-written texts with varying degrees of success (Liu et al., 2024; Mosca et al., 2023; Wang et al., 2024). Though not definitive, these results suggest there are indeed structural differences between human- and LLM-written text.
@@
In this study, we used several recent LLMs (OpenAI’s GPT-4o and GPT-4o Mini, and four variants of Meta Llama 3) to generate text from prompts drawn from a large, representative corpus of English, allowing us to directly compare the style of LLM writing to human writing. We find large differences in grammatical, lexical, and stylistic features, demonstrating that LLMs prefer specific grammatical structures and struggle to match the stylistic variation present in human writing. In Llama 3, where we are able to compare base models to instruction-tuned variants, we further see that the instruction tuning introduces more extreme grammatical differences, making them easier to distinguish from human writing and introducing features similar to those present in GPT-4o and GPT-4o Mini.
@@
For example, the instruction-tuned LLMs used present participial clauses at 2 to 5 times the rate of human text, such as in this sentence from GPT-4o using two present participles: “Bryan, leaning on his agility, dances around the ring, evading Show’s heavy blows.” They also use nominalizations at 1.5 to 2 times the rate of humans, such as in this sentence from Llama 3 70B Instruct containing four: “These schemes can help to reduce deforestation, habitat destruction, and pollution, while also promoting sustainable consumption patterns.” On the other hand, GPT-4o uses the agentless passive voice at roughly half the rate as human texts—but in each case, the Llama base models use these features at rates more closely matching humans. This suggests that instruction tuning, rather than training the models to write even more like humans, instead trains them in a particular informationally dense, noun-heavy style, and limits their ability to mimic other writing styles.
@@
These results demonstrate the value of attending to linguistic structure (morphosyntactic, functional, and rhetorical) in order to better understand the affordances and outputs of large language models. Since the rise of the Internet and the concurrent development of efficient processing architectures, language modeling has relied on relatively simple linguistic principles (i.e., sequences and context windows) as the availability of massive amounts of text allowed models trained on text to rapidly outperform older paradigms based on linguistic theory (Li, 2023); but linguistic theory can provide better ways to evaluate LLM output, just as improved benchmark problems can provide better ways to evaluate their reasoning ability. These results also demonstrate the limits of current LLMs in matching human language, showing that despite their apparent ability, they have measurable limitations compared to human authors.
2 Methods
2.1 Human-AI parallel corpora
For this study, we created two parallel corpora. The construction of each corpus was accomplished by:
  1. 1.
    assembling 12,000 human-authored texts in English representing a range of registers (or types);
  2. 2.
    selecting roughly the first 1,000 words of each text (where that span does not cross a section or chapter boundary);
  3. 3.
    splitting each string into approximately 500-word chunks (with the split-point being a sentence boundary as determined by a spaCy dependency parse);
  4. 4.
    passing the first chunk to six different LLMs and prompting the models to generate 500 more words of text in the same style, similar to the method used by Yang et al. (2024);
  5. 5.
    discarding all texts that contain a non-response or a nonsense response (like a string of punctuation) from any of the six LLMs.
The resulting corpora, then, contain the first chunk of human-authored text (the “germinal” text used for the LLM prompt) and seven variations of the second chunk: the original (human-authored) continuation of the first chunk, and continuations generated by six different LLMs. This workflow is illustrated in Figure 1. The figure also highlights the alignment of the second, parallel chunks, which enable comparisons between the writing of humans and LLMs, as well as comparisons among the writing produced by the LLMs, themselves.
@@ 2.1 Human-AI parallel corpora
Figure 1: The LLM text generation workflow. The human text was split into two chunks of roughly 500 words; the first chunk was used to prompt an LLM to create text that was compared to the second human chunk.
The first of our parallel corpora uses the Corpus of Contemporary American English (COCA) as the source of its human-authored text and is called the COCA AI Parallel (CAP) corpus. As noted above, we wanted data representative of a broad range of sources and registers. COCA is a monitor corpus with over 1 billion words and balanced among eight registers: spoken, fiction, magazines, newspapers, academic journals, blogs, other web pages, and TV/movie subtitles (Davies, 2008). For our dataset, we filtered for texts containing a minimum of approximately 1,250 words, then randomly sampled 1,500 documents from each register to create a corpus of 12,000 texts.
Though its breadth and representativeness make COCA valuable for this task, it does have a key limitation: to allow distribution of otherwise copyrighted material, COCA alters the text by removing 10 words every 200 words. [^1] See https://www.corpusdata.org/limitations.asp for details on the procedure. This produces gaps and incomplete sentences in each 500-word chunk that might affect sentence parsing and feature extraction. For this reason, we constructed a second corpus, which is meant to approximate COCA’s composition.
The resulting dataset, the Human AI Parallel English (HAP-E) corpus, is drawn from six (as opposed to COCA’s eight) text-types: [^2] The data can be accessed from https://huggingface.co/datasets/browndw/human-ai-parallel-corpus.
  • Academic: samples were drawn from a 40,000+ document corpus of open-access academic articles published by Elsevier.
  • News: samples were drawn from a 100,000+ document corpus of news articles published online by U.S.-based news organizations.
  • Fiction: samples were drawn from novels and short stories in the public domain and available on Project Gutenberg.
  • Spoken: samples were drawn from a corpus of 100,000 podcast transcriptions. (Note that COCA’s spoken text-type comes from transcriptions of unscripted news shows.)
  • Blogs: samples were drawn from a corpus of 681,288 posts from blogger.com.
  • Television and Movie Scripts: samples were drawn from two different corpora of scripts, some of which were converted using OCR.
The sampling and processing of the HAP-E corpus followed the same five steps described at the beginning of this section. The word counts of the HAP-E and CAP corpora can be found in the Appendix, Tables LABEL:tbl:totals and LABEL:tbl:cap_totals.
2.2 LLM text generation
We used multiple versions of GPT-4o (OpenAI, 2024a, b) and Meta Llama 3 (Meta, 2024) to generate the parallel LLM texts, shown in Table 1. We chose these LLMs because they are popular, available within a reasonable budget, and generally considered to perform very well. Meta Llama 3 is available in both base models, which provide text completion, and instruction-tuned variants that use additional human feedback to optimize the models to follow instructions and answer questions; recent GPT variants are only available in instruction-tuned form. Llama 3’s availability in both forms allows us to observe the effects of instruction tuning.
Model Version Context Instruct?
GPT-4o Mini 2024-07-18 128K Yes
GPT-4o 2024-08-06 128K Yes
Llama 3 8B 8K No
Llama 3 8B-Instruct 8K Yes
Llama 3 70B 8K No
Llama 3 70B-Instruct 8K Yes
Table 1: Large language models used to generate text completions.
Instruction-tuned LLMs were given the following prompt, followed by the first chunk of human text:
In the same style, tone, and diction of the following text, complete the next 500 words, generate exactly 500 words, and note that the text does not necessarily end after the generated words:
Base Llama 3 variants without instruction tuning were simply given the first chunk of text. In each case, the LLMs generated roughly 500 words of additional output, continuing where the first chunk finished. This could be compared directly to the second chunk of human text. GPT-4o text was generated using OpenAI’s Batch API using its default settings; Llama 3 text was generated on a local GPU cluster using the vLLM package for Python, version 0.5.5 (Kwon et al., 2023).
As some LLMs refused to complete some prompts, or returned short, unusable answers, we removed responses shorter than 100 words, producing 8,290 HAP-E texts and 9,615 COCA texts with valid responses from all LLMs.
We designed this procedure for creating parallel data because our aims are focused more on description than they are on detection. Compared to other methods that only prompt LLMs with a title or topic (e.g. Wang et al., 2024), this approach gives the LLMs roughly 500 words of context so they can adapt their output and its style to the human text.
2.3 Feature extraction
To extract meaningful features from our corpus for training our classifiers, we used Douglas Biber’s tagset of 66 linguistic categories, gleaned from the many works Biber has authored and co-authored (Biber, 1988, 1995; Biber and Conrad, 2009). Biber’s tagset includes indices of lexical complexity (e.g., mean word length), but many more are raw linguistic features ranging from the lexical (e.g., thinking verbs) to the grammatical (e.g., adjectives before a noun). For example, features include:
  • Nominalizations (nouns formed by adding suffixes to verbs or adjectives, such as “justification”, “development”, “robustness”)
  • Agentless passive voice (such as “The model was fitted”)
  • Hedging phrases (such as “at about”, “something like”, “almost”)
  • Phrasal coordination (such as using “and” to coordinate between noun phrases)
  • Clausal coordination (such as using “and” at the beginning of clauses).
A full list of features is provided in the Appendix, Table LABEL:tbl:all-biber-features. Combined, these variables give a stylistic sense of how these texts read. These stylistic profiles, in turn, can enrich our understandings of the affordances and limitations of LLMs as they are currently developed and trained.
To tag the texts, we used the pseudobibeR package for R, which counts Biber features present in dependency-parsed texts. [^3] Available at https://github.com/browndw/pseudobibeR pseudobibeR has proven effective at reproducing results similar to those of other studies that use Biber’s tagset (DeJeu and Brown, 2023; Markey et al., 2024). Once tagged, we counted the rate of occurrence of each feature per 1,000 words of text. These rates served as the features for our analysis.
Corpus linguists studying language use and variation have regularly turned to counts of morphosyntactic and lexical features as the basis for their quantitative analysis, an approach which dominates the field (Biber, 1988; Brezina, 2018; Hyland, 2000; Wallis, 2021). Biber’s tagset offers a compelling approach to studies of linguistic variation: because linguistic features both routinely co-occur and are sensitive to context (DeJeu and Brown, 2023), analysis of feature groupings that vary across sources lends insight into the salient distinctions between the stylistic profiles of human and LLM texts. We chose this tagset to represent this perspective, but also because it has been validated by a considerable amount of scholarship (Friginal, 2013; Hardy and Römer, 2013). Of course, other foundational corpus-based studies have used different feature sets or “bag-of-words” methods, generating word frequency lists to investigate keywords, or for instance, markers of stance in academic writing (Aull and Lancaster, 2014; Hyland, 2005), Instead, Biber’s morphosyntactic tagset presents a more robust method for analyzing register and genre, capturing a wide range of linguistic functions.
2.4 Classification
To evaluate whether the features described above are sufficient to distinguish between human- and LLM-written text, we randomly split the HAP-E documents into a training set (75%) and a test set (25%), stratifying by document category.
First, we fit a random forest to classify the source of each of the texts using the Biber features. This is a seven-class classification task (human and six LLMs). This random forest identified the features most important for distinguishing between sources; these top features will be used in Section 3 for visualizations of feature variations by source. This was fit using the ranger package (Wright and Ziegler, 2017) with its default settings.
Next, we trained a random forest for each pairwise human vs. LLM training task: human vs. GPT-4o, human vs. Llama 3 8B, and so on. For comparison, we also fit lasso-penalized logistic regression models using the same set of Biber features. We used the performance of the lasso models to benchmark and provide context to the random forest models. Since we only account for linear additive contributions of the Biber features in the lasso models, any noticeable improvements in performance of the random forests are due to the flexibility of tree-based models in capturing the presence of non-linear interactions. All lasso logistic regression models were fit using the glmnet package with the one-standard-error regularization penalty selected with 10-fold cross-validation (Friedman et al., 2010).
Finally, we used these models to classify text from the M4 parallel corpus released by Wang et al. (2024). We focus on its arXiv GPT-3.5 dataset, which includes both real abstracts of preprints posted on arXiv and text generated by GPT-3.5 when prompted to write an abstract based on the preprint’s title. This provides an external test set, generated on a different task with a different model, to evaluate the generalizability of our classifiers.
Detection has been a particular emphasis since the release of ChatGPT and similar LLMs for understandable reasons (Liu et al., 2024; Waltzer et al., 2024). While our work involves similar classification tasks, we use classification not as an end in and of itself, but rather as method for helping us understand how LLMs write: how the text that they produce compares to the writing humans produce and how the text produced by one LLM compares to that produced by another.
3 Results
3.1 Classifying text by source
@@
Figure 2: Confusion matrix for the seven-class random forest using HAP-E, evaluated on the test set. The block diagonal structure indicates that most classification errors were between different versions of the same LLM, rather than between humans and LLMs.
@@
The random forest classifier for the seven-task problem (distinguishing between human chunk 2 and the six LLMs) achieved a test accuracy of 66%, compared to an expected accuracy of 14% from random guessing. The confusion matrix, shown in Figure 2, demonstrates that little of the error was due to confusion between human texts and the LLMs: instead, most classification errors confused Llama 3 8B and 70B, 8B Instruct and 70B Instruct, or GPT-4o and 4o Mini. Each pair consists of models of two different sizes trained on similar data, implying that the size difference does not produce dramatically different style. Overall, only 4.2% of LLM texts were falsely classified as human, and only 9.8% of human texts were falsely classified as LLMs. Feature importances for all 66 features are shown in Table LABEL:tbl:top-features in the Appendix.
3.2 Differences in style and vocabulary
@@
Figure 3 illustrates the large variation in rate of occurrence of the fifteen most important features, relative to the rate observed in the human text; rates for all 66 features are provided in Table LABEL:tbl:top-features in the Appendix. All four instruction-tuned models have strong preferences for present participial clauses, ‘that’ clauses as subjects, nominalization, and phrasal co-ordination, which are typical markers of more informationally dense, noun-heavy style of writing. There are also signs of “house styles”: both GPT-4o models avoid clausal coordination, while all Llama 3 variants use it more frequently than humans; while both GPT-4o models use downtoners (such as barely or nearly) more frequently than humans, all Llama 3 variants avoid them.
@@
Figure 3: Rate of Biber feature use by different LLMs, relative to the human usage of each feature, for the top 15 most important features. Note the log scale. GPT-4o and GPT-4o Mini show the largest variation from human texts, while the base variants of Llama 3 most closely resemble human grammar and style. Larger models (top row) generally show the same rhetorical differences as their smaller counterparts (bottom row), despite performing better on other benchmark tasks.
@@
One might expect larger models to better match human text than smaller models (e.g. Llama 70B versus Llama 8B, or GPT-4o versus GPT-4o Mini), but this does not appear to be the case in Figure 3. Also, instruction tuning appears to make the model output less human, not more: the Llama 3 base models use features at rates similar to human texts, while GPT-4o and Llama 3 instruction-tuned models have much wider variation from feature to feature.
@@
Figure 4: Rates of word use by different LLMs (per 1,000 words) compared to the human use of each word in chunk 2 (log scale). Includes all words used more than once per million words in chunk 2. Words are lemmatized to group together inflected forms. Dashed blue lines indicate the range between $10\times$10 × more and $10\times$10 × less than human use. Note that the instruction-tuned models show more variation from the diagonal, indicating more deviation in vocabulary use relative to humans.
@@
Similar to past research (Liang et al., 2024), we find that LLMs also favor specific vocabulary. Figure 4 shows the rate of usage for words used more than once per million words by humans, comparing the usage of each LLM to the usage by humans in Chunk 2 of HAP-E. Compared to the base Llama models, in the instruction-tuned Llama and GPT-4o models certain words get used at dramatically higher and lower rates. For example, GPT-4o and 4o Mini use words like camaraderie, palpable, tapestry, and intricate at more than 100 times the rate of humans, such as in the GPT-4o output phrase “The camaraderie was palpable.” Conversely, they use certain obscenities more than 100 times less often. The top overrepresented and underrepresented words are presented in Tables LABEL:tbl:overrepresented and LABEL:tbl:underrepresented in the Appendix. [^4] Some overuse may be artifacts of the generation process; for example, Llama 3 instruction-tuned variants overuse continuation because their outputs sometimes begin with “Here is the continuation of the text…” Llama base models have a tendency to repeat themselves, so Llama 3 8B uses Deborah at 52 times the rate of humans largely because of a single document repeating it 308 times.
@@
While many words listed in Table LABEL:tbl:overrepresented may be occasionally expected in belletristic works of fiction, their pervasiveness across LLM output in a diverse array of genres is notable. To those familiar with academic writing, newspapers, or television scripts, these words are largely unexpected, and to experts likely signal an overwritten, sentimental, or simply uneven text. In the GPT models in particular, many of these words connote some form of complex relation among objects (e.g., tapestry, intricate, camaraderie, cacophony, amidst). Coupled with positive items such as vibrant and solace, these words together may signal a preference for grandiose, if hollow, summative sentences. As the use is most similar to humans for the base models, this suggests the word choice bias is introduced by the instruction tuning process, not simply by bias in the texts composing the training sets.
3.3 Distinguishing individual LLMs
@@
When classifying between human-generated text and one specific LLM, rather than comparing all LLMs, our classifiers achieve much higher accuracy. Typical accuracies achieved by random forests were 93–98%, as shown in Table LABEL:tbl:rf in the Appendix, even when trained on HAP-E and tested on CAP or vice versa. The lasso-penalized logistic regression classifiers attained similar performance for all LLMs except for the Llama base models, which had accuracies around 75%, as shown in Table LABEL:tbl:lasso. Since the lasso regressions only consider additive terms, this implies that interactions between the Biber features contain relevant signals for the Llama base models.

### v2

# pnasresearcharticle

@@ Abstract
Large language models (LLMs) are capable of writing grammatical text that follows instructions, answers questions, and solves problems. As they have advanced, it has become difficult to distinguish their output from human-written text. While past research has found some differences in features such as word choice and punctuation, and developed classifiers to detect LLM output, none has studied the rhetorical styles of LLMs. Using several variants of Llama 3 and GPT-4o, we construct two parallel corpora of human- and LLM-written texts from common prompts. Using Douglas Biber’s set of lexical, grammatical, and rhetorical features, we identify systematic differences between LLMs and humans and between different LLMs. These differences persist when moving from smaller models to larger ones, and are larger for instruction-tuned models than base models. This observation of differences demonstrates that despite their advanced abilities, LLMs struggle to match human stylistic variation. Attention to more advanced linguistic features can hence detect patterns in their behavior not previously recognized.
keywords:
corpus linguistics $|$ large language models $|$ natural language processing $|$ writing style
[…]
This manuscript was compiled on August 21, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
3]4
[…]
@@
As large language models (LLMs) have advanced in recent years, from “stochastic parrots” to models evidently capable of performing complex tasks, most attention has focused on their reasoning performance: solving mathematical problems, writing code, evaluating arguments, diagnosing diseases, and so on (1, 2, 3, 4). While past research has studied their mastery of basic grammar and vocabulary (5), there is relatively little research on their language performance more generally: their ability to produce readable text in a variety of styles. Rather than exploring it in detail, commentators discuss the business and communication tasks that might be automated by LLMs with their writing ability, or consider the dangers of impersonation and misinformation facilitated by LLMs (6, 7, 8, 9). As more and more writing tasks are automated, such problems appear inevitable.
@@
However, the impression that LLMs write “like humans” is based primarily on qualitative evaluation of their output, not on thorough linguistic evaluation of their text. So far, quantitative comparisons have looked mainly at basic grammar and syntax (5) or features such as word choice, punctuation, sentence length, and so on, finding evidence of some differences between human- and LLM-written text (10, 11, 12, 13, 14). Other work has used these features, or language models trained on sample texts, to classify LLM-written texts with varying degrees of success (15, 16, 17). Though not definitive, these results suggest there are indeed structural differences between human- and LLM-written text.
@@
We used several recent LLMs (OpenAI’s GPT-4o and GPT-4o Mini, and four variants of Meta Llama 3) to generate text from prompts drawn from a large, representative corpus of English, allowing us to directly compare the style of LLM writing to human writing. We find large differences in grammatical, lexical, and stylistic features, demonstrating that LLMs prefer specific grammatical structures and struggle to match the stylistic variation present in human communication, particularly as that variation aligns with the conventions that structure genres such as academic writing, interactive speech, or journalistic news. In Llama 3, where we are able to compare base models (which produce text completions) to instruction-tuned variants (which have been further trained to answer questions and complete tasks specified in prompts), we further see that the instruction tuning introduces more extreme grammatical differences, making them easier to distinguish from human writing and introducing features similar to those present in GPT-4o and GPT-4o Mini.
@@
For example, the instruction-tuned LLMs used present participial clauses at 2 to 5 times the rate of human text, such as in this sentence from GPT-4o using two present participles: “Bryan, leaning on his agility, dances around the ring, evading Show’s heavy blows.” They also use nominalizations at 1.5 to 2 times the rate of humans, such as in this sentence from Llama 3 70B Instruct containing four: “These schemes can help to reduce deforestation, habitat destruction, and pollution, while also promoting sustainable consumption patterns.” On the other hand, GPT-4o uses the agentless passive voice at roughly half the rate as human texts—but in each case, the Llama base models use these features at rates more closely matching humans. This suggests that instruction tuning, rather than training the models to write even more like humans, instead trains them in a particular informationally dense, noun-heavy style, and limits their ability to mimic other writing styles leading to, in some cases, genre misalignment.
@@
These results demonstrate the value of attending to linguistic structure (morphosyntactic, functional, and rhetorical) in order to better understand the affordances and outputs of large language models. Since the rise of the Internet and the concurrent development of efficient processing architectures, language modeling has relied on relatively simple linguistic principles (i.e., sequences and context windows) as the availability of massive amounts of text allowed models trained on text to rapidly outperform older paradigms based on linguistic theory (18); but linguistic theory can provide better ways to evaluate LLM output, just as improved benchmark problems can provide better ways to evaluate their reasoning ability. These results also demonstrate the limits of current LLMs in matching human language, showing that despite their apparent ability, they have measurable limitations compared to human authors.
Methods
@@
Figure 1: The LLM text generation workflow. Each human text was split into two chunks of roughly 500 words; the first chunk was used to prompt an LLM to create text that was compared to the second human chunk.
We created two corpora of parallel human- and LLM-written texts. Each corpus began with $n=12,000$ human-authored English texts from a range of genres, from spoken word (such as podcast transcripts) to news and magazine articles to formal academic writing. Language use varies in relation to situational factors such as audience and purpose (19, 20); by including multiple genres, we aimed to capture a diverse range of language production. As illustrated in Figure 1, from each text we extracted two consecutive chunks of roughly 500 words (split at sentence boundaries). The first chunk was provided to each LLM to give it context and a sample of the writing style. The LLMs were prompted to write 500 more words in the same style, tone, and diction; their generated text was then compared to the next 500-word chunk of the human text. We used six LLMs: GPT-4o and GPT-4o Mini (21, 22) and Meta Llama 3 8B, 70B, 8B Instruct, and 70B Instruct (23), producing six LLM-authored texts for each human-authored text. See the SI Appendix for detailed prompt information.
We constructed the first corpus, the Human-AI Parallel English corpus, from six categories of text (academic, news, fiction, spoken word, blogs, and TV/movie scripts). The second corpus, the COCA AI Parallel (CAP) Corpus, is drawn from the pre-existing Corpus of Contemporary American English (COCA), a large, representative corpus of over 1 billion words in eight registers: spoken, fiction, magazines, newspapers, academic, blogs, web pages, and TV/movie subtitles (24). The HAP-E corpus was used for our primary analyses, while CAP was used to evaluate the generalizability of the results to different texts. The LLMs sometimes refused to respond to prompts or gave short, unusable answers; after these were removed, there were $n=8,290$ HAP-E texts and $n=9,615$ CAP texts with outputs from all LLMs. With two human chunks and six LLM-authored chunks for each text, HAP-E comprised $n=66,320$ chunks and CAP $n=76,920$ chunks. See SI Appendix Tables S2–S3 for corpus size and composition.
To extract meaningful features from our corpus for training our classifiers, we used Douglas Biber’s tagset of 66 linguistic categories (19, 25, 26), which includes indices of lexical complexity and raw linguistic features ranging from the lexical to the grammatical. For example, features include mean word length, the use of nominalizations (nouns formed from adjectives or verbs, such as development or robustness), agentless passive voice, hedging phrases (such as something like or almost), and clausal coordination. All features are listed in the SI Appendix, Table S4. Differences between LLM and human use of features were tested for statistical significance with the paired Wilcoxon signed-rank test with Bonferroni multiple comparison correction.
As a further check of generalizability, we used part of the M4 parallel corpus (17) consisting of abstracts from the arXiv preprint service alongside abstracts generated by GPT-3.5 when prompted with the preprint title. These texts are from a different LLM (GPT-3.5) and a very distinct genre of writing (academic abstracts), providing a check on the consistency of results in different genres.
Results
Classifying text by source
@@ Results
Figure 2: Confusion matrix for a random forest classifying HAP-E texts by their linguistic and rhetorical features, evaluated on the test set (25% of the HAP-E corpus, including $n=14,535$ human and LLM texts). The block diagonal structure indicates that most classification errors were between different versions of the same LLM, rather than between humans and LLMs.
@@
A random forest classifier using the Biber features to distinguish between the seven text sources in HAP-E (human chunk 2 and the six LLMs) achieved a test accuracy of 66%, compared to an expected accuracy of 14% from random guessing. The confusion matrix, shown in Figure 2, demonstrates that little of the error was due to confusion between human texts and the LLMs: instead, most classification errors confused Llama 3 8B and 70B, Llama 3 8B Instruct and 70B Instruct, or GPT-4o and 4o Mini. Each pair consists of models of two different sizes trained on similar data, implying that the size difference does not produce dramatically different style. Overall, only 4.2% of LLM texts were falsely classified as human, and only 9.8% of human texts were falsely classified as LLMs.
Differences in style and vocabulary
@@
Figure 3 illustrates the large variation in rate of occurrence of the fifteen most important features (as identified by the random forest) in texts generated by LLMs, relative to the rate observed in the human text. All four instruction-tuned models have strong preferences for present participial clauses, ‘that’ clauses as subjects, nominalization, and phrasal co-ordination, which are typical markers of more informationally dense, noun-heavy style of writing (27). For example, GPT-4o uses present participial clauses at 5.3 times the rate of humans (paired Cohen’s $d=1.38$), ‘that’ clauses as subject 2.6 times as often ($d=0.77$), nominalizations 2.1 times as often ($d=1.23$), and phrasal coordination 1.9 times as often ($d=0.81$). (Rates and effect sizes for all 66 features are provided in the SI Appendix, Tables S5–S6; Figure S1 illustrates paired differences.) There are also signs of local patterns that emerge with specific models: both GPT-4o models avoid clausal coordination, while all Llama 3 variants use it more frequently than humans; while both GPT-4o models use downtoners (such as barely or nearly) more frequently than humans, all Llama 3 variants avoid them.
@@
Figure 3: Rate of Biber feature use by different LLMs, relative to the human usage of each feature, for the top 15 most important features in the HAP-E corpus. Note the log scale. GPT-4o and GPT-4o Mini show the largest variation from human texts, while the base variants of Llama 3 most closely resemble human grammar and style. Larger models (top row) generally show the same stylistic differences as their smaller counterparts (bottom row), despite performing better on other benchmark tasks. Triangles indicate statistically significant differences from human usage. Figure S1 of the SI Appendix gives the distributions of paired differences.
@@
One might expect larger models to better match human text than smaller models (e.g., Llama 70B versus Llama 8B, or GPT-4o versus GPT-4o Mini), but this does not appear to be the case in Figure 3. Also, instruction tuning appears to make the model output less human, not more: the Llama 3 base models use features at rates similar to human texts, while GPT-4o and Llama 3 instruction-tuned models have much wider variation from feature to feature.
@@
Figure 4: Rates of word use by different LLMs (per 1,000 words) compared to the human use of each word in chunk 2, in the HAP-E corpus (log scale). Includes all words used more than once per million words in chunk 2. Words are lemmatized to group together inflected forms. Words on the diagonal are used equally often in human and LLM texts. Dashed blue lines indicate the range between $10\times$ more and $10\times$ less than human use. Note that the instruction-tuned models show more variation from the diagonal, indicating more deviation in vocabulary use relative to humans.
@@
Similar to past research (14), we find that LLMs also favor specific vocabulary. Figure 4 shows the rate of usage for words used more than once per million words by humans, comparing the usage of each LLM to the usage by humans in Chunk 2 of HAP-E. Compared to the base Llama models, in the instruction-tuned Llama and GPT-4o models certain words are used at dramatically higher and lower rates. Table 1 highlights words overrepresented in LLM outputs: GPT-4o and 4o Mini use words like camaraderie, palpable, tapestry, and intricate at more than 100 times the rate of humans, such as in the GPT-4o output phrase “The camaraderie was palpable.” As a result, “tapestry” appeared in 23% of GPT-4o outputs and “amidst” in 27% (SI Appendix, Table S7). Instruction-tuned variants of Llama 3 also favor words like camaraderie and palpable, as well as unease and reminder, though at lower rates than GPT-4o and in a much smaller fraction of documents. [^1] Some overuse may be artifacts of the generation process; for example, Llama 3 instruction-tuned variants overuse continuation because their outputs sometimes begin with “Here is the continuation of the text…” Llama base models have a tendency to repeat themselves, so Llama 3 8B uses Deborah at 52 times the rate of humans largely because of a single document repeating it 308 times. Conversely, they use certain obscenities more than 100 times less often (SI Appendix Table S8).
Table 1: Most overrepresented words in LLM-generated texts, relative to human usage rates
GPT-4o GPT-4o Mini Llama 3 70B Instruct Llama 3 8B Instruct Llama 3 70B Llama 3 8B
Word Rate Word Rate Word Rate Word Rate Word Rate Word Rate
camaraderie 162 camaraderie 171 unease 63 unease 101 bananas 31 deborah 52
tapestry 155 tapestry 147 palpable 47 continuation 52 paperback 30 rambo 22
intricate 119 palpable 145 continuation 29 palpable 48 bam 26 matty 20
underscore 107 grapple 131 shoutout 28 reminder 33 verona 25 goodnight 18
unspoken 102 intricate 129 intricate 27 pang 29 filth 19 ml 15
amidst 100 fleeting 124 pang 25 rut 29 rekall 17 merlin 13
palpable 95 ignite 122 camaraderie 24 waft 28 denis 14 worcester 11
solace 95 vibrant 92 policymaker 24 prioritize 27 darry 12 fay 10
fleeting 84 amidst 90 prioritize 24 grapple 24 ebook 12 missy 10
unravel 83 cacophony 89 reminder 24 camaraderie 23 janice 12 elisa 10
@@
While many words listed in Table 1 may be occasionally expected in belletristic works of fiction, their pervasiveness across LLM output in a diverse array of genres is notable. To those familiar with academic writing, newspapers, or television scripts, these words are largely unexpected, and to experts likely signal an overwritten, sentimental, or simply uneven text. The point here is not that humans refrain from using these words, but that humans refrain from using these words in certain genres. In this case, words that are unremarkable in fiction are highly conspicuous and unconventional when used other genres. As word choice appears most similar to humans for the base models, this suggests the word choice bias is introduced by the instruction tuning process, not simply by bias in the texts composing the training sets.
In the GPT-4o models in particular, many of these words connote some form of complex relation among objects (e.g., tapestry, intricate, camaraderie, cacophony, amidst). Coupled with positive items such as vibrant and solace, these words together may signal a preference for grandiose, if hollow, summative sentences.
Distinguishing individual LLMs
@@
When classifying between human-generated text and one specific LLM, rather than comparing all LLMs, our classifiers achieve much higher accuracy. Typical accuracies achieved by random forests were 93–98% even when trained on HAP-E and tested on CAP or vice versa (SI Appendix Table S9). Lasso-penalized logistic regression classifiers attained similar performance for all LLMs except for the Llama base models, which had accuracies around 75% (SI Appendix Table S10). Since the lasso regressions only consider additive terms, this implies that interactions between the Biber features contain relevant signals for the Llama base models.